A study of 415 kindergarten students showed that they have seen on average




5000 hours of television. If the sample standard deviation of the population is




900, find the 95% confidence level of the mean for all students. If a parent




claimed that his children watched 4000 hours, would the claim be believable?

1
Expert's answer
2021-11-04T19:12:04-0400

"95\%CI=(5000-1.966\frac{900}{\sqrt{415}},5000+1.966\frac{900}{\sqrt{415}})=(4913.2,5086.8)."

The claim is not believable because 4000 does not lie in the confidence interval.